The Castle Company recently reported net profits after taxes of $14 million. It has 2 million shares of common stock outstanding and pays preferred dividends of $2 million a year. The company's stock currently trades at $159 per share.

A.- Compute the stock's earnings per share (EPS). Round the answer to two decimal places.
$ ______ per share

B.- What's the stock's P/E ratio? Round the answer to two decimal places.
$ _______ times

C.- Determine what the stock's dividend yield would be if it paid $9.38 per share to common stockholders. Round the answer to two decimal places.
_________%

If you could please show how you worked out the problem. Thank you.
Answer #1

A. EPS =Profit After Taxes/Number of Shares =14/2 =7

B. P/E Ratio =Price per Share/EPS =159/7 =22.71

C.Dividend Yield =Dividend/Price =9.38/159 =5.90%